# A Survey on Speech Large Language Models

### v2

@@ 4.2.1 Audio-to-Text Modality Conversion
    Direct Projection: The audio feature information is directly projected into the LLM’s text feature space through a connector [19, 16]. Specifically, the tensor containing audio features is projected into an embedding vector aligned with the text modality. These audio embeddings are then concatenated with the input text’s embedding vector to create a new embedding vector that integrates both speech and text information. This combined vector is then fed into the LLM.Of course, some researchers have also implicitly handled the projection step by internalizing it within the original encoder, achieving the modality projection effect by adjusting the encoder’s parameters during training [17].
@@
    Token Mapping: The audio feature information is mapped into text tokens that the LLM can process [37]. Initially, the audio features are passed through a projector or converter to generate symbols or representations that correspond to text tokens. These representations are then combined with the tokenized input text, forming a token sequence that includes both audio and text information. This sequence is subsequently input into the LLM for processing. This approach not only preserves the integrity of the audio feature information but also ensures consistent processing within the LLM.
4.2.2 Combining Audio and Text Feature Space
Projecting the speech modality into the text modality is certainly straightforward, but it does not achieve lossless modality fusion in the true sense. Information loss and conflicts may occur during modality conversion (to be elaborated later). Therefore, researchers have proposed another modality alignment approach, which involves modifying the original input space of the large language model to integrate the audio modality [14, 13, 41]. Specifically, this method augments the token space by adding audio tokens on top of the existing text tokens, creating a new token space. These new audio token are synthesized from the audio features extracted in the previous step, thereby largely preserving the original audio information (illustrated in Figure 4(c)).
5 Training Strategies
The training of current Speech LLMs mainly includes three approaches: pretraining, supervised fine-tuning (SFT), and reinforcement learning (RL). These approaches are used in different training stages and scenarios due to their varying principles and effects.
5.1 Pretraining
Pretraining involves training models on large-scale unlabeled data to enable them to learn broad, general knowledge. Typically, in Speech LLMs, speech encoders undergo pretraining on audio-text pairs to capture audio features. Common training strategies, including self-supervised learning (SSL) [18], are widely used. However, to better integrate speech encoders with LLMs, some researchers attempt to re-pretrain speech encoders [42], which allows for more direct acquisition of encoders with audio feature extraction capabilities.
For Speech LLMs of Section 4.2.2, as mentioned previously, a thorough re-training of multimodal large models is necessary. This involves referencing the pretraining phase of large language models. By organizing mixed audio and text data, we adapt the pretraining methods of large language models to accommodate bimodal information.
5.2 Supervised Fine-Tuning (SFT)
For specific tasks in the speech domain, further fine-tuning is often required. Supervised fine-tuning is a common approach where high-quality labeled data from downstream task datasets is used to train the model, improving its performance on these tasks. In Speech LLMs, for models of structure as mentioned earlier in Section 4.2.1, supervised fine-tuning is an extremely common method. It is generally used to achieve alignment between the speech encoder and the LLM and to enhance performance on specific tasks. Common training methods include fine-tuning connectors, fine-tuning the encoder, and LLMs, such as using methods like LoRA. This mainly involves handling modality alignment and completing the model’s learning of text-token mapping. Similarly, models of structure described in  Section 4.2.2 also require supervised fine-tuning, which is essential for enhancing its performance on specific tasks.
It is worth noting that supervised fine-tuning is also frequently used as a wake-up method for large language models. When modality alignment is completed, large language models might experience a certain degree of dormancy [16]. Researchers use audio-based dialogue wake-up fine-tuning to reactivate large language models.
5.3 Reinforcement Learning (RL)
Reinforcement learning (RL) is a commonly used method in the training of large language models, especially in the field of large language model safety alignment [43]. Unlike other training strategies, reinforcement learning ensures that the LLM optimizes in the desired direction while maintaining diversity in its outputs. In order to maintain and enhance cross-task integration capabilities of large language models under the Speech LLM paradigm, reinforcement learning is an effective training strategy that is worth further exploration [44, 45]. However, it has not been widely used in Speech LLMs.
Reinforcement learning is a framework where an agent interacts with an environment to maximize cumulative rewards over time. The agent learns by receiving feedback in the form of rewards, adjusting its actions to improve future rewards. Two commonly used RL algorithms in this context are Proximal Policy Optimization (PPO) and Direct Policy Optimization (DPO).
Proximal Policy Optimization (PPO) is a policy gradient method that aims to optimize the objective function while limiting policy updates to avoid large deviations from the current policy. The objective function for PPO is defined as:
$\mathcal{L}^{\text{PPO}}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (1)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ is the probability ratio between the new policy and the old policy, $\hat{A}_{t}$ is the advantage function, and $\epsilon$ is a hyperparameter that controls the clipping range. The clipping mechanism ensures that policy updates remain within a trust region, preventing overly aggressive updates that could harm the model’s performance.
Direct Policy Optimization (DPO), on the other hand, focuses on directly optimizing the policy by defining a reward function specific to the task. The objective in DPO can be expressed as:
$\mathcal{L}^{\text{DPO}}(\theta)=\hat{\mathbb{E}}_{t}\left[R(s_{t},a_{t})\log\pi_{\theta}(a_{t}|s_{t})\right]$ (2)
where $R(s_{t},a_{t})$ is the task-specific reward function, and $\pi_{\theta}(a_{t}|s_{t})$ is the policy parameterized by $\theta$. DPO directly maximizes the expected reward by focusing on optimizing the policy without clipping, often used in settings where a more direct reward signal is available.
Currently, RL is generally employed to further enhance the model’s accuracy in a specific domain by setting reward signals in a reward model to guide the model’s training [18], thereby achieving further performance improvement.
@@ 6 Performance in Speech Tasks
With the evolution of Speech LLMs, the integration of large language models (LLMs) into speech-based systems has yielded substantial advancements across various dimensions. This section investigates the current performance of Speech LLMs in several pivotal tasks in speech understanding, evaluating how LLMs enhance the capabilities of speech models compared to traditional methods. We will primarily focus on two classic speech-related tasks: Automatic Speech Recognition(ASR), and Speech Translation(ST), each epitomizing a crucial aspect of speech applications. Meanwhile, we will further explore the multi-tasking and cross-tasking capabilities exhibited by Speech LLMs across various tasks in the speech understanding field.
@@ 6.1 Automatic Speech Recognition
In this survey, we assess the performance of various models via the Word Error Rate (WER) tested on the clean and other test set of LibriSpeech dataset [46], which is an essential benchmark for evaluating ASR systems.
@@
Current Speech LLM Stage Qwen-audio 2.0 4.2
SALMONN 2.1 4.9
@@
Traditional models based on LSTM or GRU architectures have consistently demonstrated strong performance across various benchmark datasets. For instance, the LSTM-based model Deep Speech 2 achieved a Word Error Rate (WER) of approximately 3.51% on the LibriSpeech clean test set and 10.62% on the other test set [22]. The introduction of Large Language Models (LLMs) has significantly improved these results. Notably, the Conformer model reduced the WER to 1.9% and 3.9% on the clean and other test sets, respectively [8]. Similarly, the HuBERT Base model, trained solely with 960 hours of LibriSpeech audio, achieved WERs of 3.4% and 8.1%. Utilizing an extended dataset of 60,000 hours from Libri-light, the HuBERT X-Large model further reduced the WERs to 1.9% on the clean test set and 3.5% on the other test set. As of 2023, further advancements were seen with Whisper Large-V3, achieving a new low WER of 1.8% on the clean test set. Recent Speech LLM models like Qwen-audio [17] and SALMONN [16] have also shown impressive results, approaching state-of-the-art performance, with WERs around 2% on the clean test set and below 5% on the other test set. Currently, the most advanced ASR system on the LibriSpeech dataset is Seed-ASR, which achieves a WER of 1.6% on the clean test set and 2.8% on the other test set [18].
@@ 6.2 Speech Translation
In the realm of natural language processing, translation tasks, encompassing both speech translation (ST) and machine translation (MT), are pivotal for enhancing global communication. Just like other NLP fields, these translation tasks have experienced significant progress due to the recent advancements in large language models (LLMs) [47]. In this survey, we employ the BLEU (Bilingual Evaluation Understudy) score to evaluate the translation performance of each model on the two-way translation between Chinese and English. For speech translation, we utilize the FLEURS test sets [48], and for machine translation from Chinese to English, we use the FLORES test sets [49, 50]. Additionally, the WMT’20 test sets [51] are used for English-to-Chinese machine translation.
There are two primary objectives in the translation domain: enhancing translation quality and expanding language coverage. In machine translation, the NLLB model, which is fine-tuned on LLMs aiming to be a universal translation system, initially broadened its linguistic scope to include more than 200 languages [50]. Furthermore, BigTranslate, leveraging instruction tuning based on LLaMA, has expanded the translation applicability of LLMs to a broader array of languages, achieving results comparable to those of ChatGPT and Google Translate [27]. The advent of LLMs has also propelled advancements in speech translation. Using a Cascaded ASR + MT approach, Whisper has shown notable efficacy in speech translation by utilizing extensive web-scale data [23]. Subsequently, AudioPaLM2 [14] enhanced speech translation quality by integrating text and speech-based models, thereby achieving impressive results as an end-to-end speech translation model. Previously achieving state-of-the-art in both ST and MT tasks on various public datasets, SeamlessM4T [29], a foundational multilingual and multitask model, is now surpassed by GenTranslate. This model utilizes the information stored in N-best candidates instead of relying solely on beam search decoding’s top-1 selection, a common practice among other models [28]. Please refer to  Figure 6 for a detailed performance comparison of these models.
6.3 Multi-tasking and Cross-tasking Abilities
Besides the above-mentioned two speech tasks, some Speech LLMs also demonstrate remarkable results on a wide range of speech tasks, showcasing their good multi-tasking and cross-tasking capabilities. Among these, Pengi is notable for its versatility, as it supports both open-ended tasks like Audio Captioning and Audio Question Answering (AQA), and close-ended tasks such as Sound Event and Scene Classification, without requiring any task-specific fine-tuning or extensions. Pengi’s approach involves framing all audio tasks as text-generation tasks, leveraging a unified architecture that effectively blends audio and text inputs into text outputs, achieving state-of-the-art performance on several benchmarks, including audio question answering and sound event classification [15].
@@ 6.3 Multi-tasking and Cross-tasking Abilities
Lastly, Qwen-Audio significantly advances the field of audio interaction by scaling up audio-language pre-training across more than 30 tasks involving various audio types. This model excels in areas like Acoustic Scene Classification, Speech Emotion Recognition, and Automatic Audio Captioning. By employing a multi-task training framework that minimizes interference from the diversity of tasks and audio types, Qwen-Audio achieves groundbreaking results in these areas, indicating a significant enhancement in audio understanding capabilities [17].
@@ 7 Challenges
Despite significant progress in applying large language models (LLMs) to the field of speech recognition, especially in fixed tasks such as Automatic Speech Recognition (ASR), or leveraging the powerful capabilities of LLMs to accomplish some multitasking, numerous challenges and issues remain. We validated and tested some of the most cutting-edge current works, which helped to corroborate the problems we identified.
@@ 7.1 LLM’s Dormancy
Regarding the method of using a projector to project audio feature embeddings into the text token space and combining them with text token embeddings of prompts, although it shows high accuracy and considerable generalization ability in ASR tasks, the LLM does not respond positively to prompts that were not seen during training. We refer to this phenomenon as LLM dormancy. To further illustrate the issue, we use SLAM-ASR and SALMONN as replication examples:
The core idea of SLAM-ASR is projecting audio feature embeddings to text token embeddings through a projector composed of a linear neural network. Here, we use a recipe consisting of wavlm-large, a projector, and vicuna-7b, with whisper-large and vicuna-7b kept frozen during training while only the projector is trainable. We trained for three epochs on the librispeech-960h dataset and achieved results on the librispeech test_clean and test_other datasets close to the expectations outlined in the original paper.
The core idea of SALMONN is to use a window-level Q-former to integrate audio information extracted by BERT and Whisper and map it into the text token space. During training, we updated the parameters of the LLM and encoder using LoRA and the window-level Q-former. We achieved results on librispeech test_clean, test_other, and Gigaspeech [52] that matched the expectations of the original paper.
@@
We conducted a series of prompt tests on the two models mentioned above, with the results as shown in Table 2. From the table, we can see that SLAM-ASR almost does not respond to prompts, while SALMONN selectively responds to prompts, which precisely confirms our view that LLMs cannot maintain full model performance in the current framework.
Input: Can you divorce life from art?(A man speaker with chinese accent)
Output
Ability Test Prompt SLAM-LLM SALMONN
ASR transcribe speech to english text Can you divorce life from art? can you divorce life from art?
Speech Translate transcribe the speech to text and translate it to chinese. {CJK}UTF8gbsn你能从艺术中离开吗?
Speech Translate transcribe the speech to text and translate it to Japanese. {CJK}UTF8minアートから生活を破棄することができますか?
Speaker Verify which scene will this talk occur? The speaker is male and has an American accent. He speaks with confidence and clarity.
Speaker Verify Is the speaker a female? Yes, the speaker is a female.
Dialogue Ability Tell me the reason. I’m sorry, I’m not sure what you’re asking. Could you please provide more context or clarify your question?
Dialogue Ability
Based on the speech, can you help me determine
the answer?
Yes, I can divorce life from art.
Dialogue Ability can you answer the speech and give reasons? Can you divorce life from art
Context Inference
Actually this is a speech a man asked me. I don’t
have answer. And someone tells me to say yes.
But I don’t know why. Can you explain it?
Can you divorce life from art?
Yes, I can explain it. The man asked you if you could
divorce life from art. This means that you can separate
the concept of life from the artistic representation of it.
For example, a painting of a landscape could be seen
as a representation of life, but not necessarily a reflection
of it. So, the answer to the man’s question is yes, you can
divorce life from art.
Table 2: Prompt Tests conducted on Slam-ASR and SALMONN; Text highlighted in red indicates errors or incorrect outputs, while text in black represents reasonable or correct responses.
From the perspective of training strategies, SLAM-ASR did not update parameters for the encoder or LLM, thus making it more intuitive that the problem likely arises from the intermediate modality fusion step. To validate our approach, we attempted to reproduce the training process of SLAM-ASR. We then compared the average absolute value of the projected audio feature vectors with the average absolute value of the token embeddings obtained from encoding the prompt.
Fig. 7: Mean absolute value of embeddings
Note: As the tensor shape of the input embedding is (Batch size, Sequence id, Feature dimension), we set batch size to 1 and calculate the average of the absolute value of all the elements corresponding to the sequence id.
The figure here is a randomly chosen audio input for clear demonstration.
From Figure 7, we observe that the audio embeddings are significantly larger compared to the text embeddings.
Referring back to the procedure in the transformer module in a typical large language model, the text is first converted into tokens and embeddings before calculating the attention mechanism, which includes positional embeddings. However, the audio feature vectors are projected and then directly concatenated with the text token embeddings in Speech LLM. Clearly, there is no attempt to capture the positional information of the two modalities here.
A second potential issue arises during the calculation of attention scores. Recalling the standard steps involved in calculating attention scores in the attention mechanism, Given the audio-text modality embedding vector $E$, and since the LLM is kept frozen, the weight matrices $W^{Q}$, $W^{K}$, and $W^{V}$ are fixed. We transform $E$ into query, key, and value vectors using linear transformations, calculated as follows:
$Q$ $=EW^{Q}$ (3)
$K$ $=EW^{K}$ (4)
$V$ $=EW^{V}$ (5)
Next, we compute the attention scores:
Attention $\text{Attention Scores}=QK^{T}$ (6)
where $K^{T}$ is the transpose of the key vector.
To stabilize the training and inference process, we optionally scale the dot-product attention scores:
Scaled $\text{Scaled Scores}=\frac{QK^{T}}{\sqrt{d_{k}}}$ (7)
where $d_{k}$ is the dimensionality of the key vectors.
The scaled attention scores are then transformed into probability distributions using the softmax function:
[…]
Finally, we use the attention weights to perform a weighted sum of the value vectors to obtain the final attention output:
Attention Output=Attention $\text{Attention Output}=\text{Attention Weights}\cdot V$
which in detail can be expressed as:
Attention $\text{Attention Output}=\left(\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)\right)\cdot V$ (9)
Here we can observe that since the large language model is frozen, only $E$ changes based on the input. The embedding $E$ is obtained by concatenating the audio feature embedding $A$ and the text feature embedding $T$:
$E=[A\;T]$
We can also decompose $Q$, $V$, and $K$ into vectors corresponding to $A$ and $T$, for example,
$Q=[Q^{A}\;Q^{T}]$
$V=[V^{A}\;V^{T}]$
$K=[K^{A}\;K^{T}]$
Since $A\gg T$, according to equations (1), (2), and (3), due to the shared weight matrices, $Q^{A}\gg Q^{T}$ and similarly for other vectors. It is clear that the final attention output from the speech part is much larger than that from the text part. As a result, the model is likely to neglect the text portion, leading to the large language model falling into dormancy.
@@
This can also be explained more broadly and intuitively using Bayes’ theorem [16]. Let $Y$ denote the output, $X$ denote the audio feature, and $I$ denote the instruction.
$P(Y|X,I)=\frac{P(Y|X)\cdot P(I|Y,X)}{P(I|X)}$ (10)
If we provide only one or a few corresponding outputs $Y$ for a given audio input, then:
$\frac{P(Y|X,I)}{P(Y|X)}$ (11)
is not negligible. In other words, the influence of $I$ on the result will not be significant.
@@
Models like SALMONN and Qwen-audio, which are Speech LLMs, perform poorly in tasks involving untrained speech emotion inference and environmental understanding [53], and may even provide irrelevant answers. This further supports the observation that we did not fully achieve modality alignment under the current framework.

### v4

@@ 6.1.1 Manually Annotated Datasets
    Automatic Speech Recognition (ASR): LibriSpeech (English)[51], LibriLight (English)[52], GigaSpeech (English)[53], Aishell1 (Mandarin Chinese)[54], Aishell2 (Mandarin Chinese)[55], Voxpopuli (Multilingual)[56]
@@
    Speech Translation: CovoST2 (21 source languages and 15 target languages)[57], Voxpopuli (14 languages)[56]
  3. 3.
    Automatic Audio Caption: Clotho (English)[58], AudioCaps (English)[59]
Traditional multitask Datasets
In addition to single-task datasets, there are also comprehensive datasets designed to address multiple speech-understanding tasks. These datasets incorporate audio metadata besides the audio-text pair alone. For example, the AMI dataset provides rich metadata, including speaker information and discussion topics, making it applicable to a range of tasks such as ASR, Automatic Audio Captioning (AAC), ST, and SD[60]. Despite their versatility and informational depth, the relatively small size of these datasets limits their broader applicability for training Speech LLMs.
6.1.2 Datasets Annotated with LLM Participation
The rapid advancement of Large Language Models (LLM), particularly in text-based reasoning capabilities approaching human performance, has inspired researchers to explore their potential in dataset annotation. This approach typically involves representing audio information in text form and pairing it with task-specific prompts to guide LLM in generating inferences. This process, known as the SQA format (Speech [in text] + Question + Answer), allows for scalable dataset creation by reducing human labor and lowering costs. Furthermore, it provides a novel way to train LLM, facilitating general question-answering capabilities about speech that cannot be fully realized using audio-text pairs alone. During training, the text-based question and its corresponding speech context are provided as input, while the model is trained to generate answers in text form as illustrated in 5(a)
However, despite its efficiency, this method is limited by the reasoning capabilities of the LLM, which may lead to suboptimal annotations in many cases. The text prompts and answers generated by LLM often lack the diversity seen in human-created instructions. Consequently, such datasets may still fall short in supporting comprehensive speech question-answering tasks.
6.2 Datasets Generated Using TTS Technology
Given the scarcity of audio datasets specifically designed for Speech LLMs and their limited variety, researchers have proposed an innovative approach to integrate audio and text modalities more deeply to enable the model’s capability of handling flexible mixed-modal input. Building on the SQA framework, this approach converts part of the textual components (speech/questions) into speech using Text-to-Speech (TTS) technology, thereby creating Audio-Text pairs for training Speech LLMs. The structure is illustrated in 5(b).
Peng et al. introduced VoiceTextBlender, marking the first significant effort to enhance mixed-modal supervised fine-tuning (SFT) data using TTS[61]. This method allows prompts and questions to also exist in the audio modality, further enhancing the potential for multi-modal task training. This approach opens new possibilities for creating datasets that support deeper cross-modal integration while addressing the limitations of real-recorded audio datasets.
@@ 7 Performance in Speech Tasks
With the evolution of Speech LLMs, the integration of large language models (LLM) into speech-based systems has yielded substantial advancements across various dimensions. This section investigates the current performance of Speech LLMs in several pivotal tasks in speech understanding, evaluating how LLM enhance the capabilities of speech models compared to traditional methods. We will primarily focus on two classic speech-related tasks: Automatic Speech Recognition(ASR), and Speech Translation(ST), each epitomizing a crucial aspect of speech applications. Meanwhile, we will further explore the multitasking and cross-tasking capabilities exhibited by Speech LLMs across various tasks in the speech understanding.
@@ 7.1 Automatic Speech Recognition
In this survey, we assess the performance of various models via the Word Error Rate (WER) in English tested on the clean and other test set of LibriSpeech dataset [51], which is a widely recognized benchmark for Automatic Speech Recognition (ASR) systems. We also compare two leading models on a multilingual ASR task using the FLEURS test sets [62], another important benchmark for evaluating the multilingual capabilities of ASR models.
@@
Current Speech LLM Stage Firedasr-LLM 1.73 3.67
SLAM-LLM 1.84 3.39
@@
Traditional models based on LSTM or GRU architectures have consistently demonstrated strong performance across various benchmark datasets. For instance, the LSTM-based model Deep Speech 2 achieved a Word Error Rate (WER) of approximately 3.51% on the LibriSpeech clean test set and 10.62% on the other test set [63]. The introduction of Large Language Models (LLM) has significantly improved these results. Notably, the Conformer model reduced the WER to 1.9% and 3.9% on the clean and other test sets, respectively [25]. Similarly, the HuBERT Base model, trained solely with 960 hours of LibriSpeech audio, achieved WERs of 3.4% and 8.1%. Utilizing an extended dataset of 60,000 hours from Libri-light, the HuBERT X-Large model further reduced the WERs to 1.9% on the clean test set and 3.5% on the other test set. As of 2023, further advancements were seen with Whisper Large-V3, achieving a new low WER of 1.8% on the clean test set. Recent Speech LLMs models like FireRedAsr-LLEM [10] and SLAM-LLM [33] have also shown impressive results, approaching state-of-the-art performance, with WERs around 2% on the clean test set and below 5% on the other test set. Currently, the most advanced ASR system on the LibriSpeech dataset is Seed-ASR, which achieves a WER of 1.6% on the clean test set and 2.8% on the other test set [15].
To evaluate ASR performance across multiple languages, we refer to results from the Seed-ASR report, using the FLEURS dataset [62]. Compared to Whisper Large-V2, Seed-ASR (ML) demonstrates consistent improvements across all evaluated languages, including Arabic (AR), English (EN), Korean (KO), Japanese (JA), French (FR), and others. As shown in Fig. 6, Seed-ASR outperforms Whisper Large-V2 by a noticeable margin, particularly in high-error languages such as Korean and Arabic, where the gap in WER is most pronounced. This highlights Seed-ASR’s robustness and adaptability in multilingual ASR tasks, further solidifying its position as one of the most advanced ASR systems to date.
@@ 7.2 Speech Translation
In the realm of natural language processing, translation tasks, encompassing both speech translation (ST) and machine translation (MT), are pivotal for enhancing global communication. Just like other NLP fields, these translation tasks have experienced significant progress due to the recent advancements in large language models (LLM) [64]. In this survey, we employ the BLEU (Bilingual Evaluation Understudy) score to evaluate the translation performance of each model on the two-way translation between Chinese and English. For speech translation, we utilize the FLEURS test sets [62], and for machine translation from Chinese to English, we use the FLORES test sets [65, 66]. Additionally, the WMT’20 test sets [67] are used for English-to-Chinese machine translation.
There are two primary objectives in the translation domain: enhancing translation quality and expanding language coverage. In machine translation, the NLLB model, which is fine-tuned on LLM aiming to be a universal translation system, initially broadened its linguistic scope to include more than 200 languages [66]. Furthermore, BigTranslate, leveraging instruction tuning based on LLaMA, has expanded the translation applicability of LLM to a broader array of languages, achieving results comparable to those of ChatGPT and Google Translate [13]. The advent of LLM has also propelled advancements in speech translation. Using a Cascaded ASR + MT approach, Whisper has shown notable efficacy in speech translation by utilizing extensive web-scale data [18]. Subsequently, AudioPaLM2 [30] enhanced speech translation quality by integrating text and speech-based models, thereby achieving impressive results as an end-to-end speech translation model. Previously achieving state-of-the-art in both ST and MT tasks on various public datasets, SeamlessM4T [16], a foundational multilingual and multitask model, is now surpassed by GenTranslate. This model utilizes the information stored in N-best candidates instead of relying solely on beam search decoding’s top-1 selection, a common practice among other models [23]. Please refer to  Figure 7 for a detailed performance comparison of these models.
7.3 Multitasking and Cross-tasking Abilities
Besides the above-mentioned two speech tasks, some Speech LLMs also demonstrate remarkable results on a wide range of speech tasks, showcasing their good multitasking and cross-tasking capabilities. Among these, Pengi is notable for its versatility, as it supports both open-ended tasks like Audio Captioning and Audio Question Answering (AQA), and close-ended tasks such as Sound Event and Scene Classification, without requiring any task-specific fine-tuning or extensions. Pengi’s approach involves framing all audio tasks as text-generation tasks, leveraging a unified architecture that effectively blends audio and text inputs into text outputs, achieving state-of-the-art performance on several benchmarks, including audio question answering and sound event classification [31].
@@ 7.3 Multitasking and Cross-tasking Abilities
Lastly, Qwen-Audio significantly advances the field of audio interaction by scaling up audio-language pre-training across more than 30 tasks involving various audio types. This model excels in areas like Acoustic Scene Classification, Speech Emotion Recognition, and Automatic Audio Captioning. By employing a multitask training framework that minimizes interference from the diversity of tasks and audio types, Qwen-Audio achieves groundbreaking results in these areas. [19].
{CJK}UTF8gbsn Input: 甚至出现交易停滞的情况 (A woman says in Chinese, “Even there was almost no transaction” )
Output
Ability Test Prompt FireRedLLM Qwen2-Audio-Instruct
ASR
Transcribe the audio into text:
{CJK}UTF8gbsn 甚至出现交易停滞的情况
{CJK}UTF8minThe original content of this audio is:
{CJK}UTF8min’甚至出现交易停滞的情况.’
Speech Translate
Translate the speech into English:
{CJK}UTF8gbsn 甚至出现交易停滞的情况
{CJK}UTF8minThe market even几乎是停滞不前的交易情况。
Speech Translate
Translate the speech into Japanese:
{CJK}UTF8min The speech translated to Japanese is:
{CJK}UTF8min’甚至出现交易停滞的情况.’
Speaker Verify
Is the speaker a female?
Yes, the speaker is female
Dialogue Ability
Based on the speech, can you tell me
the trading situation?
The trading situation was almost stagnant.
Language Detection
What language is this voice above?
Mandarin
Emotion Recognition
What is the speaker’s emotion?
The speaker sounds frustrated.
Speaker Count
How many speakers are there?
One
Dog Detection
Is there a dog barking in the audio?
No, there is no dog barking in the audio.
Speaker Distance
How far is the speaker?
The speaker is 25 miles away.
Speaker Intent
What is the speaker’s intention?
The speaker intends to convey that there was almost a
complete stop in transactions.
Story Continuation
Continue the story
This situation is indeed worrying.
In the real estate market, a near standstill
in trading may lead to…
Table 2: Prompt Tests conducted on FireRedLLM and Qwen2-Audio-Instruct; Text highlighted in gray indicates incorrect outputs, while text in black represents reasonable or correct responses.
@@ 8 Challenges
Despite significant progress in applying LLM to the field of speech recognition, especially in fixed tasks such as ASR, or leveraging the powerful capabilities of LLM to accomplish some multitasking, numerous challenges and issues remain. We validated and tested some of the most cutting-edge current works, which helped to corroborate the problems we identified.
Fig. 8: The attention score distribution plot for the prompt ”Transcribe speech to Japanese”.
@@ 8.1 LLM’s Dormancy
Regarding the method of using a projector to project audio feature embeddings into the text token space and combining them with text token embeddings of prompts, although it shows high accuracy and considerable generalization ability in ASR tasks, the LLM does not respond positively to prompts that were not seen during training and lost their original intelligent capabilities. We refer to this phenomenon as LLM dormancy. To further illustrate the issue, we use FireRedLLM[10] and Qwen2-Audio-Instruct[12] as replication examples:
FireRedLLM is a speech LLM model focused on ASR tasks and has achieved new SOTA results across various recognition datasets. The model begins by training a Conformer architecture based on an AED framework, which serves as the audio encoder. The encoded speech tokens are then projected into the text embedding space of LLM via an projector module. The LLM is fine-tuned using LoRA to adapt its capabilities for the specific task. Notably, FireRedLLM is exclusively trained for ASR purposes, with the training prompts remaining fixed throughout the process.
Qwen2-Audio-Instruct is a multitask speech model capable of handling speech understanding tasks like speech recognition and translation, as well as conversational tasks. It uses the encoder from Whisper as its speech encoder and projects the encoded tokens into the text embedding space of LLM via a projector layer. Training is done in three stages: first, ASR and AAC tasks are trained; second, instruction fine-tuning is performed using various speech understanding tasks; and third, RLHF fine-tuning is applied to align with human conversational preferences. Notably, the training data for Qwen2-Audio-Instruct amounts to 600,000 hours.
@@
We conducted a series of prompt tests on the two models mentioned above, with the results as shown in Table 2. From the table, we can see that FireRedLLM almost does not respond to prompts, while Qwen2-Audio-Instruct incorrectly answered some questions, which precisely confirms our view that LLM cannot maintain full model performance in the current framework.
Why do speech large models lose their original intelligence? As figure 8 shows, this can be seen from the attention scores assigned to speech tokens versus text tokens. The FireRedASR model, trained solely on ASR tasks, almost entirely focuses on speech tokens. In contrast, the Qwen2-Audio-Instruct model, which uses multi-task training, shows some improvement but still exhibits a stronger preference for speech tokens compared to text tokens.
@@
This can also be explained more broadly and intuitively using Bayes’ theorem [20]. Let $Y$ denote the output, $X$ denote the audio feature, and $I$ denote the instruction. Instruction I is either completely fixed or selected from a limited set of tasks. Compared to the input audio content, instruction I lacks diversity.
$P(Y|X,I)=\frac{P(Y|X)\cdot P(I|Y,X)}{P(I|X)}\approx P(Y|X)$ (3)
@@
Other models like SALMONN and Qwen-audio also perform poorly in tasks involving untrained speech emotion inference and environmental understanding [68], and may even provide irrelevant answers. SALMONN refers to this phenomenon as task-overfit. SALMONN’s research reveals that decreasing the coefficient of Lora can enhance the diversity of model responses and mitigate task-overfit. However, this approach may lead to a decline in training task outcomes. To address this, a new activation fine-tuning stage with few shot samples was introduced. The results indicate that this method can reduce the occurrence of model task-overfit, yet the problem has not been completely resolved. This further supports the observation that we did not fully achieve modality alignment under the current framework.